# FedCC: Robust Federated Learning against Model Poisoning Attacks

### v3

@@ 4.2.3 FedCC
is an aggregation method that combines local clients’ model parameters, based on the Kernel CKA similarities between the global and local models’ PLRs. An overall algorithm is described in Appendix 0.B. A server first extracts PLRs of the global and local models and calculates the kernel CKA between them. We specifically focus on comparing PLRs because they exhibit the most distinguishable differences among differently trained models, as demonstrated in previous sections. Additionally, we utilize RBF Kernel CKA, as it effectively captures similarity differences between non-IID and malicious clients. The similarity values obtained are then clustered using a simple K-means clustering algorithm. Any clustering method that produces only two clusters can be used. Considering the assumption that the number of attackers is less than half of the total number of clients ($n/2$), we designate the clients in the larger cluster as ‘candidates.’
Since the earlier layers in CNNs learn global features and the latter layers embed local features [15, 25, 23], we used two differently normalized CKA similarity values: one normalized within the candidates (i.e., within_cka, denoted as $b$ in Figure 1) and one normalized across all the clients (i.e., across_cka, denoted as $a$ in Figure 1). For the PLR (second-last layer), we apply within_cka to the clients belonging to the larger cluster, whereas for all other layers, we apply across_cka to weighted average model parameters. By doing so, we can ensure that the end magnitude of the aggregated parameters stays similar to before the manipulations, leveraging the fact that the sum of the normalized values is one. Moreover, since our method evaluates clients in every global round and does not exclude any local models, it is also effective for adaptive attacks like DBA.
5 Experiments
5.1 Dataset and Model Architecture
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10 and CIFAR100 [14]. Details about datasets can be found in Appendix 0.A. Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN architectures for the experiments. For fMNIST, we used a 4-layer-CNN with two convolutions and two fully connected (FC) layers, incorporating dropout with probabilities of 0.25 and 0.5 before and after the FC layer, respectively. For CIFAR10, we utilized a 5-layer-CNN with three convolutions, each followed by a max-pool layer and two FC layers, applying dropout with a probability of 0.5 before the first FC layer. For CIFAR100, we adopted LeNet [16].
5.2 Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use Dirichlet distribution, which is widely used and well-known for its capability of representing real-world data [18]. A concentration parameter $\alpha$ controls the degree of non-IID; the smaller $\alpha$ is, the more likely the clients hold examples from only one randomly chosen class. We set the $\alpha$ to 0.2 to represent non-IID, based on [11, 38].
5.3 Experimental Setup and Baselines
We implement two untargeted attacks from [7] and two targeted attacks from [3] and [33], along with defense baselines including Krum, multi-Krum [4], Coomed [34], Bulyan [10], FLTrust [6], FLARE [30], and FedCC. We use ten clients with a 1.0 client participation fraction unless stated otherwise. As in [30], we have three adversaries for each untargeted attack and one for the targeted attack. In the DBA, we randomly select four malicious clients and implant five distributed triggers per batch. To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied. Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data. Benign clients undergo three local epochs, while compromised clients undergo six, with training done using the Adam optimizer (learning rate = 0.001).
5.4 Evaluation Metrics
6 Results and Discussion
In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments. We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [33]. Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading the detection by anomaly detection. We show the result of CIFAR10, with various defense methods applied.
6.1 Non-IID Data Environment
Table 2: Test Accuracy under untargeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum non-IID fM 64.63 16.51 45.91 57.12 13.39 49.37 49.54 66.27
C10 33.69 15.38 20.09 35.70 19.23 36.63 17.03 37.40
C100 2.27 1.00 5.34 4.85 0.98 11.62 7.46 14.51
Fang -Med non-IID fM 16.32 49.33 66.30 68.90 64.12 12.38 52.25 72.76
C10 10.02 25.06 40.52 40.23 32.47 12.39 14.59 42.37
C100 1.00 6.24 13.97 10.27 6.91 1.00 1.00 16.12
6.1.1 Untargeted Model Poisoning Attacks in non-IID setting
Table 2 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment. A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients. Under Fang-Krum, Krum, Bulyan, and multi-Krum experience significant accuracy drops (16.51%, 13.39%, and 45.91%, respectively, compared to FedAvg’s 64.63%), reflecting the attack’s design to degrade Krum-based defenses. Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings. FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates. FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients. In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates. FedCC demonstrates the highest accuracy (66.27%, 37.40%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values. Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively. FLARE continues to perform poorly for a similar reason mentioned earlier. Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers. Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments. Finally, FedCC achieves the highest accuracy (72.76%, 42.37%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance. This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
6.1.2 Targeted Backdoor Attacks
Table 3: Test Accuracy under targeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
Table 3 summarizes test accuracy under targeted attacks in a non-IID setting. To further illustrate the impact, Figure 3 provides a visual representation of backdoor confidence. It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients. In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy. FLARE reduces main task accuracy for reasons similar to those observed in previous experiments. FedCC achieves the highest test accuracy. Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training. This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median. A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation.
Notably, Figure 3 demonstrates that FedCC significantly reduces backdoor confidence. Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity. Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions. This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions. While other baselines, like Bulyan or FLTrust, also reduce confidence, we argue that FedCC not only preserves but potentially increases accuracy by precisely distinguishing malicious clients.
6.2 IID Data Environment
Table 4: Test Accuracy under untargeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
6.2.1 Untargeted Model Poisoning Attacks
6.2.2 Targeted Backdoor Attacks
Table 5: Test Accuracy under targeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks. A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy. In contrast, the fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter out outliers effectively. Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and distributed backdoor attacks (DBA), consistently neutralizing the backdoor task. These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
6.3 Performance in Varying FL Settings
6.3.1 Impact of the Number of Malicious Clients
Figure 5: Test Accuracy with different numbers of malicious clients
6.3.2 Impact of Fraction
Untargeted-Krum Untargeted-Med
Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
7 Conclusion and Future Work
FL has emerged in response to the rising concerns about privacy breaches while employing AI techniques. FL, however, is vulnerable to model poisoning attacks due to a server’s blindness to local datasets, making it difficult to assume data distribution and model parameter integrity. In line with this, many robust aggregation algorithms and defenses have been proposed; still, existing approaches are orthogonal to various attacks and lack consideration of non-IIDness. Throughout the exhaustive experiments, FedCC mitigates both untargeted and targeted (or backdoor) attacks while demonstrating its effectiveness in non-IID data environments. We leave theoretical guarantees of FedCC and experiments when clients have different neural network architectures or tasks as future work.
Appendix
Appendix 0.A Dataset and Model Architecture
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [32], CIFAR10, and CIFAR100 [14].
fMNIST consists of 60,000 training and 10,000 testing samples of 28$\times$28 grayscale images and associating labels of 10 classes. The labels include T-shirt/top, Trouser, Pullover, Dress, Coat, Sandal, Shirt, Sneaker, Bag, and Ankle boot.
CIFAR10 consists of 50,000 training and 10,000 testing examples of 32$\times$32 color images and associating labels of 10 classes. The classes include airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
CIFAR100 is the same as CIFAR10 except for the categorization. It is labeled as 100 classes, each containing 600 images. Thus, there are 500 training and 100 testing images per class. Each class (i.e., ‘fine’ label) belongs to one of the 20 superclasses.
Appendix 0.B FedCC: Algorithm
Algorithm 1 summarizes FedCC’s workflow.
@@ Appendix 0.B FedCC: Algorithm
      Output: agg_w, larger_cluster_members
$\triangleright$ Get PLRs of local models

### v4

@@ 4.2.3 FedCC
is an aggregation method that combines local clients’ model parameters based on the Kernel CKA similarities between the global and local models’ PLRs. The complete algorithm is provided in Algorithm 1. The server first extracts the PLRs from both the global and local models, then computes the RBF Kernel CKA between them. We focus on PLRs specifically, as they exhibit the most distinguishable differences between benign and malicious models, as shown in earlier sections. The RBF Kernel CKA is chosen for its effectiveness in capturing nuanced similarity differences under non-IID and adversarial settings. The resulting similarity values are clustered into two groups using a simple K-means algorithm—though any binary clustering method would suffice. Assuming that fewer than half of the clients are adversarial ($n/2$), the larger cluster is designated as the set of ‘candidates.’ This approach does not require manual thresholding; the fixed $K=2$ clustering aligns with the practical assumption that fewer than half of the clients are adversarial, and empirically, this dynamic separation has shown consistent effectiveness across rounds and datasets.
Given that earlier CNN layers capture global features while later layers encode local ones [16, 26, 24], we apply two types of normalization: within_cka (denoted $b$ in Figure 2), computed within the dominant (larger) cluster, and across_cka (denoted $a$), computed across all clients. For the PLR (second-to-last layer), we use within_cka; for all other layers, we use across_cka to compute layer-wise weighted averages. This helps preserve the scale of the aggregated parameters, as the weights are normalized to sum to one.
FedCC is also designed to remain effective against adaptive attacks, including distributed adaptive attacks like DBA. Unlike defenses based on outlier detection in parameter space, FedCC evaluates representational similarity using Kernel CKA between PLRs of local and global models. This enables the detection of semantic deviations even when malicious updates mimic benign gradients. By clustering clients based on CKA similarity and applying soft, layer-specific weighting—rather than hard rejection—FedCC attenuates the influence of suspicious clients without excluding them outright. This makes it robust against attacks that aim to blend in with benign behavior, particularly under non-IID conditions where traditional defenses often fail.
Theoretical Insight.
While FedCC is primarily supported by empirical evidence, we briefly offer a theoretical perspective on its robustness. Let $\mathcal{R}_{i}$ denote the PLR of client $i$, where benign clients’ PLRs are drawn from distribution $\mathcal{D}_{b}$ and malicious clients’ from $\mathcal{D}_{a}$. If the inter-client Kernel CKA similarity within $\mathcal{D}_{b}$ is significantly higher than between $\mathcal{D}_{b}$ and $\mathcal{D}_{a}$, then clustering based on PLR similarity can effectively separate benign from malicious clients. Prior work has empirically demonstrated that PLRs from benign clients exhibit high intra-distribution similarity across diverse data distributions [31]. FedCC leverages this separability by softly weighting updates: clients close to the majority cluster receive higher aggregation weights, while those that deviate receive less influence. This approach mitigates adversarial updates without relying on hard rejection, and remains effective even when malicious updates mimic benign gradients. Therefore, under the mild assumption that benign clients maintain high intra-cluster PLR similarity and form the majority, FedCC is theoretically expected to preserve robustness across a range of attack scenarios.
@@
      Output: agg_w, larger_cluster_members      $\triangleright$ Get PLRs of local models
@@
5 Experiments
5.1 Dataset and Model Architecture
We use three benchmark vision datasets: Fashion-MNIST (fMNIST) [33], CIFAR10, and CIFAR100 [15]. Considering that the end devices are normally incapable of handling heavy computation due to resource or communication constraints, we used lightweight CNN models for experiments. Specifically, we use 4-layer, 5-layer CNN for fMNIST and CIFAR10, and LeNet [17] for CIFAR100, as summarized in Table 2.
Table 2: CNN Architectures for fMNIST, CIFAR-10, and CIFAR-100
Dataset Layer In Out Ker / Str / Pad Activation
fMNIST conv2d_1 1 64 $5\times 5$ / 1 / 0 ReLU
conv2d_2 64 64 $5\times 5$ / 1 / 0 ReLU
dropout - - 0.25 -
flatten - 25600 - -
fc_1 25600 128 - -
dropout - - 0.5 -
fc_2 128 10 - -
CIFAR-10 conv2d_1 3 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d - - $2\times 2$ / - / 0 -
conv2d_2 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d - - $2\times 2$ / - / 0 -
conv2d_3 64 64 $3\times 3$ / 1 / 0 ReLU
maxpool2d - - $2\times 2$ / - / 0 -
flatten - 256 - -
dropout - - 0.5 -
fc_1 256 128 - -
fc_2 128 10 - -
CIFAR-100 conv2d_1 3 6 $5\times 5$ / 1 / 0 ReLU
maxpool2d - - $2\times 2$ / 2 / 0 -
conv2d_2 6 16 $5\times 5$ / 1 / 2 ReLU
maxpool2d - - $2\times 2$ / 2 / 0 -
flatten - $4$
fc_1 44944 120 - ReLU
fc_2 120 84 - ReLU
fc_3 84 100 - -
5.2 Non-IID Simulation
Non-IID is a more feasible assumption considering the diverse and massive nature of data in practice. To standardize non-IID settings, we use the Dirichlet distribution, which is widely adopted for modeling client-level label skew [19]. The concentration parameter $\alpha$ controls the degree of heterogeneity; smaller values lead to clients holding examples from fewer classes. We set $\alpha=0.2$ to simulate moderate heterogeneity, following prior works [12, 38, 31]. This value is commonly used as it balances realism and trainability: smaller values (e.g., $\alpha<0.1$) can induce excessive label imbalance and hinder convergence, while larger values (e.g., $\alpha>0.5$) reduce the heterogeneity to near-IID conditions [5].
5.3 Experimental Setup and Baselines
We implement two untargeted attacks from [8] and two targeted attacks from [3] and [34], along with defense baselines including Krum, multi-Krum [4], Coomed [35], Bulyan [11], FLTrust [7], FLARE [31], and FedCC. We use ten clients with a 1.0 client participation fraction unless stated otherwise. As in [31], we have three adversaries for each untargeted attack and one for the targeted attack. In DBA, we randomly select four malicious clients and implant five distributed triggers per batch. To compensate for delayed attack effects, we use FedAvg until epoch 30, after which each defense method is applied. Note that, for a fair comparison, FLTrust is computed using its previous global model without auxiliary data. Benign clients undergo three local epochs, while compromised clients undergo six, with training done using Adam optimizer (learning rate = 0.001).
5.4 Evaluation Metrics
6 Results and Discussions
In this section, we report the performance of our FedCC against two untargeted and two targeted attacks in both IID and non-IID environments. We additionally demonstrate the defense effect against a distributed adaptive attack, DBA [34]. Unlike centralized backdoor attacks, the DBA splits a trigger pattern across multiple adversaries, evading detection by anomaly detection. We show the result of CIFAR10, with various defense methods applied.
6.1 Non-IID Data Environment
6.1.1 Untargeted Model Poisoning Attacks in Non-IID Setting
Table 3 provides the global model’s accuracy across three datasets under two untargeted (Fang-Krum and Fang-Med) in a non-IID environment. A notable finding is that robust aggregation algorithms often yield lower accuracy than simple averaging (fedAvg). This discrepancy arises from the imperfect identification of malicious clients, leading to the erroneous aggregation of their weights along with those of benign clients. Under Fang-Krum, Krum, multi-Krum, and Bulyan experience significant accuracy drops (16.51%, 45.88%, and 13.39%, respectively, compared to FedAvg’s 57.14%), reflecting the attack’s design to degrade Krum-based defenses. Coomed also struggles, as median weights fail to represent benign clients effectively in non-IID settings. FLTrust mitigates Fang-Krum but performs better on CIFAR10 than fMNIST, likely due to fMNIST’s simpler gradients and lower variance, which challenge its ability to distinguish malicious updates. FLARE’s lower accuracy stems from its approach of scaling the entire model weights uniformly, which ignores knowledge contributed by individual clients. In contrast, FedCC, which selectively averages weights layer-wise, achieves superior performance by boosting or minimizing weights based on their alignment with benign updates. FedCC demonstrates the highest accuracy (71.13%, 52.06%, and 14.51% for fMNIST, CIFAR10, and CIFAR100, respectively) and shows significant improvement in CIFAR100 experiments.
Turning to the Fang-Med attack, FedAvg experiences a significant drop in accuracy due to the attack’s creation of outliers by deviating from median values. Consequently, Coomed and Bulyan perform relatively well since their coordinate-wise median and trimmed mean methods disregard outliers effectively. FLARE continues to perform poorly for a similar reason mentioned earlier. Between the Krum-based methods, multi-Krum and Coomed are more robust than Krum, as they aggregate multiple local models, providing better resilience to outliers. Similarly, Bulyan’s averaging of multiple clients enhances robustness but falls short of Coomed, particularly in non-IID environments. Finally, FedCC achieves the highest accuracy (72.76%, 47.85%, and 16.12% for each dataset) without sharing raw data, showcasing its exceptional performance. This success is attributed to two factors: (1) highly distinguishable information within PLRs and (2) higher CKA similarity between benign clients than between benign and malicious clients.
Table 3: Test Accuracy under untargeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum non-IID fM 57.14 16.51 45.88 57.12 13.39 60.8 49.54 71.13
C10 33.69 15.38 20.5 35.7 19.23 41.87 17.03 52.06
C100 2.27 1 4.95 7.85 0.98 11.04 7.46 14.51
Fang -Med non-IID fM 16.32 49.33 66.84 68.9 64.12 18.96 52.25 72.76
C10 10.02 25.06 45.44 40.23 32.47 10 14.59 47.85
C100 1 6.24 14.52 10.27 6.91 1.09 1 16.12
6.1.2 Targeted Backdoor Attacks in Non-IID Setting
Table 4 summarizes test accuracy under targeted attacks in a non-IID setting. To further illustrate the impact, Figure 4 provides a visual representation of backdoor confidence. It is evident that Krum, which selects a single client’s weights as the global model, shows reduced robustness compared to methods that aggregate weights from multiple clients. In contrast, Coomed, multi-Krum, and Bulyan exhibit higher test accuracy.
FLARE reduces main task accuracy for reasons similar to those observed in previous experiments. FedCC achieves the highest test accuracy. Since targeted attacks aim to misclassify a specific target class while maintaining the main task accuracy, the best performance of FedCC is attributed to leveraging knowledge from earlier stages of training. This result highlights the superiority of using sophisticated weighted knowledge and filtering malicious clients via CKA similarity, which outperforms first-order statistical methods like mean or median.
A similar trend is observed with the DBA attack, except for multi-Krum; the lowest accuracy occurs because all four malicious clients, each with a distributed trigger, are treated as outliers and excluded from aggregation. In contrast, FedCC preserves high main task accuracy (52.28%) even under DBA, outperforming all baselines. This result highlights its robustness against adaptive strategies that distribute triggers across multiple compromised clients—a setting specifically designed to evade anomaly-based filtering.
Table 4: Test Accuracy under targeted attacks in Non-IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target non-IID fM 75.65 45.27 65.97 71.70 57.96 61.82 64.31 75.66
C10 36.16 14.98 30.72 48.97 40.11 44.06 10.18 51.56
C100 4.46 6.18 6.90 12.04 11.16 12.95 1.14 15.26
DBA C10 38.56 24.94 7.09 44.45 34.19 51.49 38.73 52.28
Notably, Figure 4 demonstrates that FedCC significantly reduces backdoor confidence. Note that since the backdoor confidence of DBA fluctuates a lot due to its distributed nature, we omit it for brevity. Unlike other methods, such as FLTrust or multi-Krum, which exhibit fluctuating confidence values, FedCC maintains consistently low confidence, emphasizing its resilience and independence from client-specific data distributions. This advantage stems from the effective utilization of CKA, enabling similarity calculations even when models are trained on different datasets with varying distributions. While other baselines, like Bulyan or FLTrust, also reduce confidence, FedCC not only mitigates the backdoor attack but also sustains—or improves—accuracy by precisely identifying and down-weighting malicious clients.
Summary of Non-IID Robustness.
Across all attack types, FedCC consistently demonstrates superior performance in non-IID environments—outperforming prior defenses in both main task accuracy and backdoor confidence. This robustness stems from the use of Kernel CKA over penultimate layer representations, which captures distributional structure even under client heterogeneity. Unlike methods based on geometric or statistical outlier filtering, FedCC’s representation-based, layer-aware weighting enables it to preserve benign signals while attenuating subtle or distributed malicious updates, making it particularly effective under realistic non-IID conditions.
6.2 IID Data Environment
6.2.1 Untargeted Model Poisoning Attacks in IID Setting
Table 5: Test Accuracy under untargeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Fang -Krum IID fM 75.55 31.66 87.78 87.62 50.30 89.53 79.16 89.57
C10 49.67 40.86 63.42 57.40 12.67 68.25 25.77 69.84
C100 13.72 1.04 7.64 6.17 1.59 17.14 7.49 18.47
Fang -Med IID fM 20.86 85.33 89.53 86.70 87.45 21.36 71.08 89.66
C10 9.51 54.28 69.68 59.20 57.69 9.92 49.82 70.52
C100 0.87 12.27 16.52 14.43 12.56 1.16 5.83 17.83
6.2.2 Targeted Backdoor Attacks in IID Setting
Reducing backdoor confidence is as critical as maintaining main task accuracy, as attackers aim to stealthily embed backdoor tasks. A key observation is FedCC’s remarkable ability to reduce backdoor confidence to near zero while preserving high main task accuracy. In contrast, fluctuating backdoor confidence observed in Multi-Krum and Flare is largely due to the distributed trigger, which confuses their defense mechanisms and hampers their ability to filter outliers effectively. Other methods, such as Multi-Krum, Flare, and FLTrust, show significant variability in backdoor confidence, further highlighting FedCC’s superiority.
Notably, FedCC achieved the highest main task accuracy across both targeted backdoor attacks and DBA, consistently neutralizing the backdoor task. These results underscore FedCC’s ability to mitigate targeted backdoor attacks through precise and efficient integration of prior knowledge to preserve main task performance while zeroing out the attack confidence.
Table 6: Test Accuracy under targeted attacks in IID setting.
Case data FedAvg Krum MKrum Coomed Bulyan FLTrust FLARE FedCC
Target IID fM 88.27 86.63 87.03 89.41 89.45 89.59 75.29 90.01
C10 64.68 57.69 71.19 69.85 68.76 68.61 11.09 71.64
C100 13.83 5.72 17.57 13.24 15.11 17.08 1.03 18.61
DBA C10 10.00 35.58 51.40 10.00 16.16 56.69 10.00 58.04
6.3 Performance in Varying FL Settings
6.3.1 Impact of the Number of Malicious Clients
Figure 6: Test Accuracy with different numbers of malicious clients
6.3.2 Impact of Fraction
Untargeted-Krum Untargeted-Med
Frac Data FedAvg Med FLARE FedCC FedAvg Med FLARE FedCC
0.1 fM 55.31 49.83 34.02 64.83 16.57 66.41 52.24 69.52
C10 10.06 22.55 14.50 20.49 10.00 15.33 10.00 29.81
0.3 fM 64.22 57.52 10.00 73.55 16.26 58.07 10.00 61.12
C10 24.24 12.59 10.00 27.81 10.98 22.61 10.00 38.27
0.5 fM 62.37 58.20 10.00 76.41 18.36 62.49 10.00 69.05
C10 23.99 17.28 10.06 34.32 9.87 27.83 10.00 37.27
7 Limitation
While FedCC demonstrates strong performance against various poisoning attacks in both IID and non-IID settings, several limitations remain.
First, our experiments focus on lightweight CNN architectures (e.g., LeNet and custom CNNs) and relatively small-scale vision datasets such as fMNIST, CIFAR10, and CIFAR100. While these choices reflect practical FL deployments on resource-constrained devices, the generalizability of FedCC to larger-scale datasets (e.g., ImageNet) or more complex architectures (e.g., ResNet, ViTs) remains untested. Additionally, FedCC currently assumes homogeneous model architectures across clients. Extending it to heterogeneous or multi-task client settings is an important direction for future work.
Second, while FedCC does not rely on explicit access to raw data and has shown strong empirical robustness, including against adaptive attacks like DBA, it lacks formal guarantees under broader adversarial conditions. Theoretical analysis of its robustness bounds, especially under targeted mimicry of benign PLR distributions, remains open. Our theoretical insight provides a preliminary foundation but invites further exploration.
Third, while Kernel CKA offers strong representational comparison performance, it incurs computational overhead during aggregation. Although this cost is manageable in our experimental setting, optimizing its computation or exploring more efficient alternatives would be necessary for scalability in large-scale deployments.
8 Conclusion and Future Work
FL has emerged in response to growing concerns about data privacy in collaborative machine learning. However, the distributed nature of FL introduces vulnerabilities to model poisoning attacks, especially under non-IID data settings. While many existing defenses are effective under specific threat models, they often neglect the challenges introduced by client heterogeneity.
In this paper, we proposed FedCC, a robust aggregation method that leverages Kernel CKA to measure representational similarity in the penultimate layer of client models. Through extensive experiments across multiple datasets and attack types, we demonstrated that FedCC effectively mitigates both untargeted and targeted (backdoor) attacks, even under severe non-IID conditions.
For future work, we plan to provide theoretical guarantees for FedCC’s robustness and extend it to scenarios with heterogeneous model architectures. We also aim to evaluate its generalizability on larger and more complex datasets (e.g., ImageNet) and explore its effectiveness against stealthier or adaptive attack strategies. Lastly, we will investigate optimization techniques to reduce the computational cost of Kernel CKA, enabling scalable deployment in large-scale federated systems.